# thanks: Corresponding author

### v1

@@ III Influence of the Reference Sample on the Derived qMFM Tip Transfer Function
Figure 3: Comparison of spectral densities for different-sized reference patterns (left column) and the tip field distribution $\mu_{0}H_{z}^{\text{tip}}$ measured by qMFM magnetometry. The right column compares the $\mu_{0}H_{z}^{\text{tip}}$ obtained from NV calibration using relevant reference patterns with the $\mu_{0}H_{z}^{\text{tip}}$ measured by qMFM magnetometry. In (a), the spectral density of the CoPt sample’s reference pattern, as measured with MFM, is compared to the spectral density of the magnetic tip measured by qMFM magnetometry. The $\mu_{0}H_{z}^{\text{tip}}$ plotted in purple represents the field distribution obtained from a forward simulation using the original reference pattern as a control. From top to bottom, domain sizes of the reference patterns are increased by factors of 2 (b), 3 (c), and 4 (d). The blue, orange, and green plots correspond to different pixel resolutions and image sizes, see Table 1.
@@
The left panel of Fig. 3 shows the Fourier spectral density of the artificially generated reference patterns with different parameters, along with the Fourier spectrum of $B_{z}^{\text{NV}}$ (grey) and the Fourier spectrum of the CoPt multilayer reference sample (light red). Note that the spectra are plotted against wavelengths $\lambda$ in microns (feature size) instead of wave vector. The right panel shows the real-space representation of $B_{z}^{\text{NV}}$ (red) and the experimentally derived $B_{z}^{\text{qMFM}}$ (black), along with the real-space representations of TTFs extracted from simulated reference measurements using various reference patterns. Additionally, the right panel of Fig. 3a shows the TTF calculated from a simulated measurement of the original reference sample as described (purple). This initial simulation serves as the control. The good agreement between measured and simulated TTFs confirms that the underlying model describes the MFM calibration process reasonably well, indicating the validity of our forward simulations for differently-sized reference patterns.
@@
The spectral density plots of the original reference pattern and the tip field distribution $B_{z}^{\text{NV}}$ exhibit only a small overlapping area. This overlap (or the lack thereof) should limit the spectral information that can be reconstructed in any qMFM calibration. Specifically, a narrow Fourier spectrum of the reference sample should limit the Fourier components of the tip stray field distributions that are accessible through the calibration. Such limitations constrain both the validity, range, and quality of the reconstructed tip stray field distribution.
@@
To validate this hypothesis, we compare results from a standard CoPt reference sample (Fig. 3a) with those obtained from TTFs derived via simulated calibrations using the artificially generated reference samples as described above. These samples (see Table 1) exhibit larger domains, varied scan image sizes, and different resolutions, thereby altering their Fourier spectra compared to the original CoPt patterns. The colors in the table correspond to those in the plots.
@@
In Fig. 3b to d, the blue curve represents data with a smaller image size, while the orange and green curves correspond to larger image sizes. Among these, the orange plot displays higher resolution (i.e., smaller pixel size) compared to the green curve. Generally, the average feature size of the reference patterns increases from Fig. 3b to d. Note that the orange and green curves overlap almost exactly in all right-hand side panels. They differ only in the pixel size (i.e. image resolution) of the simulated measurement of the reference sample, suggesting that within the considered frequency range, resolution is of minor relevance for reconstructing the stray field distribution.
@@
As seen in the insets in the left panel of Fig. 3a, the feature size of the reference sample increases from a to d. Consequently, in the left-hand side panels, the spectrum’s maximum for the orange/green curves and of the blue curve shifts to the right and thus a larger value of $\lambda$. Simultaneously, as $\lambda$ increases the TTF depicted on the right panels increasingly resembles the NV-measured field distribution. This trend supports our assumption: as the overlapping area between the reference sample’s Fourier spectrum and the tip stray field distribution expands, the reconstructed TTF more closely aligns with the NV-measured tip field (red line).
@@
To determine if our findings are reproducible under experimental conditions, we compare TTFs derived from calibration measurements of two distinct reference samples using different MFM tips. These samples include the previously discussed CoPt reference sample and a TiPtCo multilayer stack detailed in Ref. Sakar et al., 2021a. Both exhibit maze domain patterns but with differing characteristic domain widths, as illustrated in Fig. 4a and b.
@@
The domain width of the TiPtCo stack is $D_{\text{TiPtCo}}=345\,\text{nm}$, whereas that of the CoPt stack is $D_{\text{CoPt}}=235\,\text{nm}$. These widths were calculated using a self-correlation approach (see Ref. Sakar et al., 2021a). Consequently, this results in different Fourier spectra, depicted in Fig. 4c, where the peaks of the TiPtCo magnetic surface charge distribution are shifted to higher wavelengths $\lambda$.
It is important to note that these samples not only differ in domain width but also in thickness and saturation magnetization, leading to a lower stray field amplitude for the TiPtCo sample.
@@
Figure 4: Comparison of TTFs derived from calibrations based on MFM measurements of two different reference samples. The CoPt reference sample (a) and the TiPtCo reference sample (b) exhibit distinct domain pattern widths, which are also reflected in their circularly averaged Fourier spectra (c). The TTF derived from the TiPtCo sample is notably wider than that obtained from the CoPt sample (d).
[…]
@@
IV Influence of the Derived qMFM Tip Transfer Function on the Reconstructed Stray Field Distribution in qMFM
@@
So far, we have analyzed how different Fourier spectra in MFM calibration patterns affect the resulting TTFs. In this section, we explore the impact of these derived TTFs on qMFM measurements of stray field distributions for typical magnetic nanostructures, namely a skyrmion-like sample, and an artificial QR code-like magnetic pattern. The ICFs are calculated for the TTFs obtained from the simulated reference patterns. By deconvolving the simulated MFM test data with these ICFs, we derive surface charge density patterns and field distributions of the test simulations. These results are then compared with the fields calculated from the sample magnetization used in the simulation.
[…]
@@
The simulated MFM phase shift data were then deconvolved using different ICFs derived from the TTF data obtained for various artificial reference samples, as discussed in the previous section. We compared these fields with those calculated directly from the magnetization patterns used to simulate the test samples.
@@
Figure 5: Artificial skyrmion sample. The top row shows the two-dimensional $\mu_{0}H_{z}$ distributions of skyrmions obtained by simulation (a), deconvolution using the transfer function from the original reference pattern (b), and from the largest reference pattern (x4, c). The field in (d) is deconvolved using the distribution obtained from NV measurements. In (e), the plot lines representing the maximum values of the skyrmion field distributions are compared. The inset figure provides the transfer functions used for these deconvolutions.
@@
The most accurate result is obtained from deconvolution using the experimental TTF (Fig. 5b and Fig. 5e – black). Conversely, the stray field distribution derived from $\mu_{0}H_{z}^{\text{tip}}$ calibrated using a reference pattern four times larger is the least accurate (Fig. 5c and Fig. 5e – red). This discrepancy arises due to the absence of higher frequency $k$-vectors in the Fourier spectrum of the reference pattern.
@@
The field distribution deconvolved using $\mu_{0}H_{z}^{\text{tip}}$ measured by NV is similarly poor (Fig. 5d and Fig. 5 – orange). This suggests that the interpolation of $\mu_{0}H_{z}^{\text{tip}}$, achieved by fitting NV data with a multipole model, failed to accurately reconstruct higher frequency $k$-vectors in the Fourier spectrum. These vectors are predominantly induced by magnetization in the tip-apex area. We assume, that the 100 nm grid used in the NV measurement leads to a loss of information. Furthermore, the significant distance between the tip and NV during the measurement of $B_{z}^{\text{NV}}$ likely resulted in a loss of contributions from the rapidly decaying high-frequency components of the tip stray field distribution.
Figure 6: Artificial QR code sample: In the top row, the two-dimensional $\mu_{0}H_{z}$ of the magnetic structure is shown, obtained by simulation (a), deconvolution using the transfer function from the original reference pattern (b) (note that this dataset is scaled by a factor of 0.2), the reference pattern with doubled domain size (x2, c), and the largest reference pattern (x4, d). The field in (e) is derived from deconvolution using the field distribution obtained from NV measurements. All plots, except for the qMFM data (factor 0.2), share the same color scale. In (f), plot lines representing the maximum values of the skyrmion field distributions are compared. The inset figure shows the transfer functions used for these deconvolutions.
@@
Figure 6 presents the derived stray field data $\mu_{0}H_{z}$ for the QR code pattern. The worst result, which significantly deviates from the simulated stray field data, arises from deconvolution using the experimental TTF (Fig. 6b and Fig. 6f – black). Note, that in this figure, the derived amplitudes for this dataset had to be divided by a factor of 5 (marked x 0.2 in Fig. 6b) to plot them alongside the other results. In addition to this strong deviation of the field amplitude, the spatial structure of the simulated stray field is not well resolved and captures only a few basic features. The $\mu_{0}H_{z}^{\text{tip}}$ data reconstructed from TTFs derived from larger reference patterns (x2 and x4) are much more accurate (Figures 6c, 6d, and 6f – green, red) which can be attributed to the higher density of $k$-vectors in the relevant spatial frequency range. Note that the best reconstruction results from deconvolution using the $B_{z}^{\text{NV}}$ data (Fig. 6e and Fig. 6f – orange).
@@
Figure 7 compares the spectra of the reference sample with our two test samples. Previously, the Fourier spectra were presented as a function of wavelength (shown in the inset of this figure); however, they are now plotted in the main figure as a function of wave vector to better resolve relevant features. The Fourier spectrum of the reference sample (green shading) shows good overlap with that of the skyrmion sample (blue) but lacks features at low wavelengths that are essential for accurately representing the QR code (red), which explains the observed results.
@@
V Summary and Conclusion
In summary, we calculated a set of tip transfer functions (TTFs) from simulated magnetic force microscopy (MFM) measurements of artificial reference samples with different characteristic feature sizes. These were compared to the TTF derived from a CoPt reference sample measurement and the tip stray field data obtained through quantum calibration using an NV center-based measurement. The simulations showed good agreement with experimental data, demonstrating the reliability of the approach.
Both the simulated and measured TTFs based on the original pattern of the CoPt reference sample exhibit deficiencies in low-wavelength distributions, resulting in significant deviations from the $B_{z}^{\text{NV}}$ data. On the other hand, while the NV-based measurement offers a quantum-based evaluation of the tip’s stray field distribution, it lacks high-frequency contributions likely due to the interpolation model employed and the sparse density of measurement points near the peak of the stray field distribution. We assume that information about the stray field induced by magnetization near the tip apex is compromised.
TTFs are crucial for quantitative MFM measurements as they enable the calculation of the instrument calibration function (ICF), which is essential for deconvolving MFM data to determine effective magnetic surface charge density or the stray field distribution of a SUT. Our findings indicate that an accurate reconstruction of a SUT’s measured stray field distribution can be achieved with a TTF, provided there is substantial overlap between the Fourier spectra of the SUT and the reference sample used for determining the TTF. A wide spectral overlap is required to achieve a high measurement signal above the noise level for the relevant spectral components. Remarkably, even if a TTF significantly deviates from the actual tip stray field distribution, it can still yield very precise reconstructions, provided there is sufficient overlap. Therefore, ensuring a proper match between the Fourier spectra of the reference sample and the SUT is critical for obtaining reliable quantitative MFM (qMFM) data. Reference samples featuring maze domain patterns are particularly well-suited for analyzing magnetic structures with characteristic feature sizes on the 100 nm scale. For larger feature sizes, new reference samples with micrometer-scale features need to be developed, tested, and validated. This will extend qMFM capabilities from the nanometer to the micrometer scale and help bridge the calibration gap to techniques with lower spatial resolution like magneto-optical indicator film measurements Dorosinskiy and Sievers (2023).

### v2

@@ IV Influence of the Reference Sample on the Derived qMFM Tip Transfer Function
Figure 4: Comparison of spectral densities (left column) for different-sized reference patterns (middle column). The right column compares the $\mu_{0}H_{z}^{\text{tip}}$ obtained from qMFM calibration using relevant reference patterns with the $\mu_{0}H_{z}^{\text{tip}}$ measured by NV magnetometry. In (a), the spectral density of the CoPt sample’s reference pattern, as measured with MFM, is compared to the spectral density of the magnetic tip measured by NV magnetometry. The $\mu_{0}H_{z}^{\text{tip}}$ plotted in purple represents the field distribution obtained from a forward simulation using the original reference pattern as a control. From top to bottom, domain sizes of the reference patterns are increased by factors of 2 (b), 3 (c), and 4 (d). The blue, orange, and green plots correspond to different pixel resolutions and image sizes, see Table 1.
@@
The left panel of Fig. 4 shows the Fourier spectral density of the artificially generated reference patterns with different parameters, along with the Fourier spectrum of $B_{z}^{\text{NV}}$ (grey) and the Fourier spectrum of the CoPt multilayer reference sample (light red). Note that the spectra are plotted against wavelengths $\lambda$ in microns (feature size) instead of wave vector. The middle panels display the simulated reference patterns at a fixed pixel size of 10 nm and a resolution of 512 px.
The right panel shows the real-space representation of $B_{z}^{\text{NV}}$ (colored red) and the experimentally derived $B_{z}^{\text{qMFM}}$ (colored black), along with the real-space representations of TTFs extracted from simulated reference measurements using various reference patterns. Additionally, the right panel of Fig. 4a shows the TTF calculated from a simulated measurement of the original reference sample as described (purple). This initial simulation serves as the control. The good agreement between measured and simulated TTFs confirms that the underlying model describes the MFM calibration process reasonably well, indicating the validity of our forward simulations for differently-sized reference patterns.
@@
The spectral density plots of the original reference pattern and the tip field distribution $B_{z}^{\text{NV}}$ in Fig. 4a exhibit only a small overlapping area. This overlap (or the lack thereof) should limit the spectral information that can be reconstructed in any qMFM calibration. Specifically, a narrow Fourier spectrum of the reference sample should limit the Fourier components of the tip stray field distributions that are accessible through the calibration. Such limitations constrain both the validity, range, and quality of the reconstructed tip stray field distribution.
@@
To validate this hypothesis, we compare results from a standard CoPt reference sample (Fig. 4a) with those obtained from TTFs derived via simulated calibrations using the artificially generated reference samples as described in Section III. These simulated scans (see Table 1) exhibit larger domains, varied scan image sizes, and different resolutions, thereby altering their Fourier spectra compared to the original CoPt patterns.
@@
In Fig. 4b to d, the colors used in the figure correspond to those listed in Table 1. Consequently, the blue curve represents data with a smaller image size, while the orange and green curves correspond to larger image sizes. Among these, the orange plot displays higher resolution (i.e., smaller pixel size) compared to the green curve. Generally, the average feature size of the reference patterns increases from Fig. 4b to d. Note that the orange and green curves overlap almost exactly in all right-hand side panels. They differ only in the pixel size (i.e. image resolution) of the simulated measurement of the reference sample, suggesting that within the considered frequency range, resolution is of minor relevance for reconstructing the stray field distribution.
@@
As seen in the insets in the middle column of Fig. 4, the feature size of the reference sample increases from a to d. Consequently, the spectrum’s maximum for the orange/green curves and of the blue curve shifts to the right and thus a larger value of $\lambda$. Simultaneously, as $\lambda$ increases the TTF depicted on the right panels increasingly resembles the NV-measured field distribution. This trend supports our assumption: as the overlapping area between the reference sample’s Fourier spectrum and the tip stray field distribution expands, the reconstructed TTF more closely aligns with the NV-measured tip field (colored red).
@@
To determine if our findings are reproducible under experimental conditions, we compare TTFs derived from calibration measurements of two distinct reference samples. These samples include the previously discussed CoPt reference sample and a TiPtCo multilayer stack detailed in Ref. Sakar2021. Both exhibit maze domain patterns but with differing characteristic domain widths, as illustrated in Fig. 5a and b.
@@
The domain width of the TiPtCo stack is $D_{\text{TiPtCo}}=345\,\text{nm}$, whereas the domain width of the CoPt stack is $D_{\text{CoPt}}=235\,\text{nm}$. These widths were calculated using a self-correlation approach (see Ref. Sakar2021). Consequently, this results in different Fourier spectra, depicted in Fig. 5c, where the peaks of the TiPtCo magnetic surface charge distribution are shifted to higher wavelengths $\lambda$. This Fourier spectrum was calculated by Fourier-transforming the two-dimensional measurements shown in Fig. 5a and b, followed by circular averaging of the resulting Fourier images. Thus, all Fourier components in all directions are considered.
@@
Figure 5: Comparison of TTFs derived from calibrations based on MFM measurements of two different reference samples. The CoPt reference sample (a) and the TiPtCo reference sample (b) exhibit distinct domain pattern widths, which are also reflected in their circularly averaged Fourier spectra (c). The TTF (here visualized by a cross-section) derived from the TiPtCo sample is notably wider than that obtained from the CoPt sample (d).
The TTFs were calculated from these measurements as previously discussed. Given the different Fourier spectra of the samples, it was anticipated that using the same tip field would yield a somewhat wider TTF in real space after calibration with the TiPtCo multilayer stack. This expectation is confirmed by the cross sections (which are taken in the x-direction through the center of the resulting TTFs, analogous to those in Fig.2) shown in Fig.5d. The comparison reveals a wider TTF for calibrations using the TiPtCo sample, which has a broader domain width. This effect was not evident in Ref. Sakar2021, where stray field gradients were analyzed, which are less sensitive to higher $k$-components due to their slower decay behavior and consequently lower gradients. The change in the TTF is less pronounced than the differences observed in Fig. 5. We primarily attribute this to a lower ratio of domain widths, of only $D_{\text{TiPtCo}}/D_{\text{CoPt}}=1.47$, and the use of a different tip, here. Furthermore, the lower stray field amplitude of the TiPtCo sample, combined with a low moment tip, results in a reduced signal-to-noise (S/N) ratio. Despite of these differences, we regard the results as at least a qualitative confirmation of our simulation-based findings.
It is important to note that these samples not only differ in domain width but also in thickness and saturation magnetization, leading to a lower stray field amplitude for the TiPtCo sample. In our experience, the CoPt sample can withstand medium moment tips without exhibiting tip-sample interaction; however, this is not the case for the TiPtCo sample, which leads to domain movement and an uneven distribution of dark and bright domains. To avoid modifications to the TiPtCo sample during imaging, a Nanosensors SSS-QMFMR low-moment tip was used for this measurement. Both samples were subsequently measured using the same tip in a Park Systems NX-Hivac MFM under ambient conditions at a lift height of 60 nm. This class of low-moment tips has previously been verified to not cause tip-sample interaction with the CoPt sample Hu2020. The TiPtCo sample has been demonstrated to withstand a MFM_LM tip Sakar2021, which has a higher moment than the SSS-QMFMR tip, making the latter a safe choice for this measurement.
@@
V Influence of the Derived qMFM Tip Transfer Function on the Reconstructed Stray Field Distribution in qMFM
@@
So far, we have analyzed how different Fourier spectra in MFM calibration patterns affect the resulting TTFs. In this section, we explore the impact of these derived TTFs on qMFM measurements of stray field distributions for typical magnetic nanostructures. The ICFs are calculated for the TTFs obtained from the simulated reference patterns. By deconvolving the simulated MFM test data with these ICFs, we derive surface charge density patterns and field distributions of the test simulations. These results are then compared with the fields calculated from the sample magnetization used in the simulation.
The two “typical” test samples represent structures with different characteristic length scales and are shown in Fig. 6a and 7a, respectively. Fig. 6a shows an artificial skyrmion lattice with a skyrmion diameter of 100 nm, whereas Fig. 7a shows a micrometer-scale QR code-like structure. The skyrmion lattice was generated using the Romming formula Romming2015. For the QR code, a purely perpendicular magnetization configuration with alternating up and down magnetized areas (light/dark shades) was assumed. For both simulated samples, parameters included a film thickness of $t=25\,\text{nm}$ and a saturation magnetization of $M_{\text{S}}=1.4\times 10^{5}\,\text{A/m}$. From these magnetization distributions, the effective magnetic surface charge density and the sample’s stray field at any distance were calculated.
The simulated MFM phase shift measurement data were generated by convolving the effective magnetic surface charge density with the ICF calculated from the NV-measured $B_{z}^{\text{NV}}$ distribution, which represents the “real” tip stray field distribution. A cantilever stiffness of $C=3.063\,\text{N/m}$, and a Q-factor of the cantilever of $Q=160.7$ were used. The stray field was calculated at a distance of 80 nm from the sample surface.
@@
The simulated MFM phase shift data were then deconvolved using different ICFs derived from the TTF data obtained for various artificial reference samples, as discussed in Section IV. We compared these fields with those calculated directly from the magnetization patterns used to simulate the test samples.
@@
Figure 6: Artificial skyrmion sample. The left column shows the two-dimensional $\mu_{0}H_{z}$ distributions of skyrmions obtained by simulation (a), deconvolution using the transfer function from the original reference pattern (b), and from the largest reference pattern (x4, c). The field in (d) is deconvolved using the distribution obtained from NV measurements. In (e), the plot lines representing the maximum values of the skyrmion field distributions are compared. The inset figure provides the transfer functions used for these deconvolutions.
Artificial Skyrmion Sample
@@
The most accurate result is obtained from deconvolution using the experimental TTF (Fig. 6b and Fig. 6e – black). Conversely, the stray field distribution derived from $\mu_{0}H_{z}^{\text{tip}}$ using a reference pattern four times larger is the least accurate (Fig. 6c and Fig. 6e – green). This discrepancy arises due to the absence of higher frequency $k$-vectors in the Fourier spectrum of the reference pattern.
@@
The field distribution deconvolved using $\mu_{0}H_{z}^{\text{tip}}$ measured by NV is similarly poor (Fig. 6d and Fig. 6 – red). This suggests that the interpolation of $\mu_{0}H_{z}^{\text{tip}}$, achieved by fitting NV data with a multipole model, failed to accurately reconstruct higher frequency $k$-vectors in the Fourier spectrum. These vectors are predominantly induced by magnetization in the tip-apex area. We assume, that the 100 nm grid used in the NV measurement leads to a loss of information. Furthermore, the significant distance between the tip and NV during the measurement of $B_{z}^{\text{NV}}$ likely resulted in a loss of contributions from the rapidly decaying high-frequency components of the tip stray field distribution.
Artificial QR Code Magnetic Pattern
@@
Figure 7 presents the derived stray field data $\mu_{0}H_{z}$ for the QR code pattern. The worst result, which significantly deviates from the simulated stray field data, arises from deconvolution using the experimental TTF (Fig. 7b and Fig. 7f – black). Note, that in this figure, the derived amplitudes for this dataset had to be scaled by a factor of 0.2 (marked “/5” in Fig. 7b) to plot them alongside the other results. In addition to this strong deviation of the field amplitude, the spatial structure of the simulated stray field is not well resolved and captures only a few basic features.
The $\mu_{0}H_{z}^{\text{tip}}$ data reconstructed from TTFs derived from larger reference patterns ($\times 2$ and $\times 4$) are much more accurate (Figures 7c, 7d, and 7f – blue and green) which can be attributed to the higher density of $k$-vectors in the relevant spatial frequency range. Note that the best reconstruction results from deconvolution using the $B_{z}^{\text{NV}}$ data (Fig. 7e and Fig. 7f – red).
The significantly poorer performance of the QR code sample (compared to the artificial skyrmion sample) when deconvolving the sample via the experimental TTF can be explained by examining the Fourier spectra of both SUTs and the reference sample.
Figure 7: Artificial QR code sample: In the top row, the two-dimensional $\mu_{0}H_{z}$ of the magnetic structure is shown, obtained by simulation (a), deconvolution using the transfer function from the original reference pattern (b) (note that this dataset is scaled by a factor of 0.2), the reference pattern with doubled domain size (x2, c), and the largest reference pattern (x4, d). The field in (e) is derived from deconvolution using the field distribution obtained from NV measurements. All plots, except for the qMFM data (factor 0.2), share the same color scale. In (f), plot lines (along the dashed line) representing the maximum values of the skyrmion field distributions are compared. The inset figure shows the transfer functions used for these deconvolutions.
Fourier Spectra of Skyrmion-, QR-, and Reference-Sample
@@
Fig. 8 compares the spectra of the reference sample with our two test samples. Previously, the Fourier spectra were presented as a function of wavelength (shown in the inset of this figure); however, they are now plotted in the main figure as a function of wave vector to better resolve relevant features. The Fourier spectrum of the reference sample (green shading) shows good overlap with that of the skyrmion sample (blue) but lacks features at low wavelengths that are essential for accurately representing the QR code (red), which explains the observed results.
@@
VI Summary and Conclusion
We calculated a set of tip transfer functions (TTFs) from simulated magnetic force microscopy (MFM) measurements of artificial reference samples with different characteristic feature sizes. These were compared to the TTF derived from a CoPt reference sample measurement and the tip stray field data obtained through quantum calibration using an NV center-based measurement. The simulations showed good agreement with experimental data, demonstrating the reliability of the approach.
However, both the simulated and measured TTFs based on the original pattern of the CoPt reference sample exhibit deficiencies in low-wavelength distributions, resulting in significant deviations from the $B_{z}^{\text{NV}}$ data. Moreover, while the NV-based measurement offers a quantum-based evaluation of the tip’s stray field distribution, it lacks high-frequency contributions likely due to the interpolation model employed and the sparse density of measurement points near the peak of the stray field distribution. We assume that information about the stray field induced by magnetization near the tip apex is compromised.
TTFs are crucial for quantitative MFM measurements as they enable the calculation of the instrument calibration function (ICF), which is essential for deconvolving MFM data to determine the stray field distribution of a sample under test (SUT). Our findings indicate that an accurate reconstruction of a SUT’s measured stray field distribution can be achieved if substantial overlap between the Fourier spectra of the SUT and the reference sample used for determining the TTF is ensured. Remarkably, even if a TTF significantly deviates from the actual tip stray field distribution, it can still yield very precise reconstructions, provided there is sufficient overlap.
Consequently, reference samples featuring maze domain patterns are particularly well-suited for analyzing magnetic structures with characteristic feature sizes on the 100 nm scale. For larger feature sizes, new reference samples with micrometer-scale features need to be developed, tested, and validated. This will extend qMFM capabilities from the nanometer to the micrometer scale and help bridge the calibration gap to techniques with lower spatial resolution like magneto-optical indicator film measurements Dorosinskiy2023.